# Dependency Parsing is More Parameter-Efficient with Normalization

### v1

@@ Abstract
Dependency parsing is the task of inferring natural language structure, often approached by modeling word interactions via attention through biaffine scoring. This mechanism works like self-attention in Transformers, where scores are calculated for every pair of words in a sentence. However, unlike Transformer attention, biaffine scoring does not use normalization prior to taking the softmax of the scores. In this paper, we provide theoretical evidence and empirical results revealing that a lack of normalization necessarily results in overparameterized parser models, where the extra parameters compensate for the sharp softmax outputs produced by high variance inputs to the biaffine scoring function. We argue that biaffine scoring can be made substantially more efficient by performing score normalization. We conduct experiments on six datasets for semantic and syntactic dependency parsing using a one-hop parser. We train $N$-layer stacked BiLSTMs and evaluate the parser’s performance with and without normalizing biaffine scores. Normalizing allows us to achieve state-of-the-art performance with fewer samples and trainable parameters. Code: https://anonymous.4open.science/r/EfficientSDP-70C1
@@ 1 Introduction
Dependency parsing (DP) consists in classifying node labels $t_{i}$ (words), edges $e_{ij}$ (relations), and edge labels $r_{ij}$ (relation types) of a dependency graph qi2019universal . A popular model for this task, introduced by Dozat and Manning dozat_deep_2017 , entails modeling word interactions as a fully connected graph via biaffine attention. Despite its simplicity, a number of models using this biaffine transformation dozat_deep_2017 ; dozat_simpler_2018 ; ji_graph-based_2019 ; jiang_entity-relation_2024 ; bhatt_end–end_2024 require more parameters than necessary, due to what we identify as a lack of normalization of its outputs. We propose that this overparameterization is caused by the high variance of the outputs, which extra parameters help mitigate. After showing that variance can be reduced through normalization, we demonstrate that similar or better performance can be obtained for DP with fewer parameters and training samples.
In this task, we consider a graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$ comprising a set of nodes $\mathcal{V}$, connected by a set of edges $\mathcal{E}$, with each edge $e_{ij}\in\mathcal{E}$ connecting pairs of nodes $(v_{i},v_{j})\in\mathcal{V}$. In latent graph inference for dependency graphs, a sentence of $|\mathcal{V}|$ words is modeled as a $|\mathcal{V}|\times|\mathcal{V}|$ graph via biaffine scoring dozat_deep_2017 :
@@
Observe that this is equivalent to the unnormalized scores used in vaswani_attention_2017 ’s self-attention: $QK^{\top}/\sqrt{d_{k}}\$ where $\text{Attention}(Q,K,V)=\text{Softmax}\left(QK^{\top}/\sqrt{d_{k}}\right)V$ and $d_{k}$ is the output size of the keys.
The main difference resides in the fact that Transformer attention scores are scaled by $a=1/\sqrt{d_{k}}$. As explained in vaswani_attention_2017 , this scaling is done because high variance inputs to the softmax function will result in large values dominating the output ($e^{x\gg 1}$) and small values decaying ($e^{x\ll 1}$). The downstream effect is exploding and vanishing gradients. To see how this works, observe that the score $s$ between any query and key vector is the result of their dot product $q_{i}k_{i}^{\top}$. Now if we assume that these vectors are zero mean and unit variance, then the variance is $\text{Var}(s)=d_{k}$. Finally, we get our normalization factor, which ensures that each entry in the score matrix has a standard deviation of 1, since $\text{Std}=\sqrt{d_{k}}$. Consequently, lower input variance will result in more stable outputs.
We observe that there is no consistency in the literature on the use of this scaling term for DP. Most works do not use this scaling term  dozat_deep_2017 ; dozat_simpler_2018 ; ji_graph-based_2019 ; bhatt_end–end_2024 ; dozat_simpler_2018 ; jiang_entity-relation_2024 , with the exception of tang-etal-2022-unirel who use vaswani_attention_2017 ’s self-attention out of the box. In this paper, we carry out a series of experiments on semantic and syntactic DP tasks, using a wide range of architecture ablations. We extend the work of Bhatt et al. bhatt_end–end_2024 , which represents the state of the art on dependency graph parsing in NLP.
@@
The rest of the paper is structured as follows. Section 2 presents an overview of the literature concerning DP, with specific focus on how to infer the adjacency matrix of a dependency graph. Section 3 explains why layer depth can compensate for normalization. Section 4 provides an overview of the datasets, models, evaluation, and other experimental details. Section 5 illustrates how normalization mitigates overparameterization. Finally, Section 6 provides avenues for future research, such as applications of our work to more complicated graph neural network-based approaches.
@@ 2 Background
In general, DP is an NLP task concerned with inferring the dependency graph of a sentence qi2019universal . Depending on the nature of the nodes and relations of the graph, it can e.g., take on the name of semantic (SemDP) dozat_simpler_2018 ; tang-etal-2022-unirel ; bhatt_end–end_2024 ; jiang_entity-relation_2024 or syntactic (SynDP) dozat_deep_2017 ; yang2018scidtb ; ji_graph-based_2019 . In this work, we tackle both tasks, but focus more specifically on SemDP, which can be thought of as comprising two sub-tasks: Named Entity Recognition (NER) nadeau_survey_2009 and Relation Extraction (RE) bhatt_end–end_2024 ; jiang_entity-relation_2024 . They can be either approached in a pipeline as separate objectives chan2011exploiting ; yan-etal-2021-partition ; zelenko2003kernel ; zhong-chen-2021-frustratingly , or by jointly training a single model on both sub-tasks bhatt_end–end_2024 ; bi_codekgc_2024 ; jiang_entity-relation_2024 ; zaratiana_autoregressive_2024 . In the context of deep learning models trained end-to-end, three main paradigms are used: encoder-based models bhatt_end–end_2024 ; dozat_deep_2017 ; dozat_simpler_2018 ; ji_graph-based_2019 ; jiang_entity-relation_2024 ; wadden-etal-2019-entity ; wang-lu-2020-two , decoder-only Transformers (LLMs) bi_codekgc_2024 ; wei_chatie_2024 ; zhang_extract_2024 , and seq2seq encoder-decoder Transformers liu-etal-2022-autoregressive ; lu-etal-2022-unified ; paolini2021structured ; zaratiana_autoregressive_2024 .
In this work, we focus on encoder-style models, since they currently achieve the best performance on DP tasks tang-etal-2022-unirel ; jiang_entity-relation_2024 ; bhatt_end–end_2024 and are much more parameter efficient than LLM-based solutions. These models approach entity (node) prediction analogously to NER, while edges and relations are handled via MLP projection of node-pair features kiperwasser_simple_2016 ; pei_effective_2015 ; peng-etal-2017-deep ; zhu-etal-2019-graph , or via attention-based edge and relation inference bhatt_end–end_2024 ; dozat_simpler_2018 ; ji_graph-based_2019 ; jiang_entity-relation_2024 ; tang-etal-2022-unirel . Current state-of-the-art models, exemplified by bhatt_end–end_2024 , are based on the biaffine dependency parser introduced by dozat_deep_2017 , which uses stacked BiLSTM layers on top of embeddings from a pre-trained language model.
@@ 3 Layer depth can compensate for normalization
We explain this discrepancy using insights from the theory of implicit regularization arora_implicit_2019 , which states that, as layer depth increases, the effective rank of the weight matrices is reduced during gradient descent. Our claim is that a reduction in the rank of the weight matrices causes a corresponding decrease in the variance of the input to the softmax function.
@@
Minimization of a loss function $\mathcal{L}(W)$ with gradient descent using weight matrices $\mathbf{W}$ (assuming a small learning rate $\eta$ and initialization close to the origin). An $N$-layer linear neural network leads the singular values of  $\mathbf{W}$ to evolve in the number of iterations $t$ by:
$\sigma_{r}(t+1)\leftarrow\sigma_{r}{t}-\eta\cdot\langle\nabla\mathcal{L}(W(t)),\mathbf{u}_{r}(t)\mathbf{v_{r}}^{\top}(t)\rangle\cdot N\cdot\sigma_{r}(t)^{2-2/N}$
This implies that, as the number of layers increases, the rank of the weight matrices decreases, since the smallest singular values decay.
Figure 1: Effective rank $\rho(W)$ reduces over training epochs as we increase $N$ of BiLSTM layers.
While Result 1 is stated for deep linear neural networks for the matrix factorization task, it has been validated empirically as applying to deep non-linear neural networks razin2020implicit ; zhao2022combining . We validate this finding for the biaffine scoring setting in Figure 1, which shows the inverse proportionality between the effective rank $\rho(W)$ roy2007effective of a BiLSTM’s weights and the number of its layers.
Claim 1 (Rank variance inequality).
Let $Y_{r}\in\mathbb{R}^{m}$ be a random vector that is the outcome of a linear transformation $\mathbf{y}=\mathbf{A}_{r}\mathbf{x}$, where the vector $\mathbf{x}\in\mathbb{R}^{n}$ is drawn from the random vector $X$ and $\mathbf{A}_{r}\in\mathbb{R}^{m\times n}$ denotes a rank $r$ approximation of a fixed matrix $\mathbf{A}$. The variance of $Y_{r}$ is smaller than the variance of $Y_{r+1}$ as measured by the Frobenius norm $||\text{Var}(Y_{r})||_{F}\leq||\text{Var}(Y_{r+1})||_{F}$.
@@
The variance of a random vector $X$ is represented by the positive semi-definite (PSD) covariance matrix $\mathbf{K_{xx}}\in\mathbb{R}^{n\times n}$. Applying the variance of $X$ to the linear transformation, we get $\text{Var}(Y)=\mathbf{A_{r}}\mathbf{K_{xx}}\mathbf{A_{r}}^{\top}\in\mathbb{R}^{m\times m}$.
Since $\mathbf{K_{xx}}$ is PSD, so is each term in $\mathbf{x_{i}^{\top}K_{xx}x_{i}}$, and so the matrix $\mathbf{A_{r}}\mathbf{K_{xx}}\mathbf{A_{r}}^{\top}$ is also PSD for each rank-$r$ approximation $\mathbf{A}_{r}=\sum_{i=1}^{r}\sigma_{i}\mathbf{u}_{i}\mathbf{v}_{i}^{\top}$.
This means that increasing the rank from $r-1$ to $r$ adds a new PSD component $\Delta_{r}$:
$\mathbf{A_{r}}\mathbf{K_{xx}}\mathbf{A_{r}}^{\top}=\mathbf{A_{r-1}}\mathbf{K_{xx}}\mathbf{A_{r-1}}^{\top}+\Delta_{r}$
Since the Frobenius norm is given by $\sqrt{\sum^{m}_{i}\sum_{j}^{n}|a_{ij}|^{2}}$, then we must have:
$||\mathbf{A_{r}}\mathbf{K_{xx}}\mathbf{A_{r}}^{\top}||_{F}$ $=||\mathbf{A_{r-1}}\mathbf{K_{xx}}\mathbf{A_{r-1}}^{\top}+\Delta_{r}||_{F}$
$\geq||\mathbf{A_{r-1}}\mathbf{K_{xx}}\mathbf{A_{r-1}}^{\top}||_{F}$
$=||\mathbf{A_{r-2}}\mathbf{K_{xx}}\mathbf{A_{r-2}}^{\top}+\Delta_{r-1}||_{F}$
$\geq||\mathbf{A_{1}}\mathbf{K_{xx}}\mathbf{A_{1}}^{\top}||_{F}$
Which shows that, as the rank decreases, so does the variance of the linear transformation.
@@
Table 1: Samples per partition and entity/relation classes for the datasets used in this paper.
Data (Train / Dev / Test) Entities Relations
ADE 2,563 / 854 / 300 gurulingappa2012ade disease, drug adverseEffect
CoNLL04 922 / 231 / 288 roth2004conll04 organization, person, location kill, locatedIn, workFor, orgBasedIn, liveIn
SciERC 1,366 / 187 / 397 luan2018scierc generic, material, method, metric, otherSciTerm, task usedFor, featureOf, hyponymOf, evaluateFor, partOf, compare, conjunction
ERFGC 242 / 29 / 29 yamakata_english_2020 food, tool, duration, quantity, actionByChef, discontAction, actionByFood, actionByTool, foodState, toolState agent, target, indirectObject, toolComplement, foodComplement, foodEq, foodPartOf, foodSet, toolEq, toolPartOf, actionEq, timingHeadVerb, other
enEWT (UD 2.2) 10,098 / 1,431 / 1,427 yamakata_english_2020 xPOS tags UD relations
SciDTB 2,567 / 814 / 817 yang2018scidtb xPOS tags UD relations

### v2

@@ Abstract
Dependency parsing is the task of inferring natural language structure, often approached by modeling word interactions via attention through biaffine scoring. This mechanism works like self-attention in Transformers, where scores are calculated for every pair of words in a sentence. However, unlike Transformer attention, biaffine scoring does not use normalization prior to taking the softmax of the scores. In this paper, we provide theoretical evidence and empirical results revealing that a lack of normalization necessarily results in overparameterized parser models, where the extra parameters compensate for the sharp softmax outputs produced by high variance inputs to the biaffine scoring function. We argue that biaffine scoring can be made substantially more efficient by performing score normalization. We conduct experiments on semantic and syntactic dependency parsing in multiple languages, along with latent graph inference on non-linguistic data, using various settings of a $k$-hop parser. We train $N$-layer stacked BiLSTMs and evaluate the parser’s performance with and without normalizing biaffine scores. Normalizing allows us to achieve state-of-the-art performance with fewer samples and trainable parameters. Code: https://github.com/paolo-gajo/EfficientSDP
@@ 1 Introduction
Dependency parsing (DP) consists in classifying node labels $t_{i}$ (words), edges $e_{ij}$ (relations), and edge labels $r_{ij}$ (relation types) of a dependency graph qi2019universal . A popular model for this task, introduced by Dozat and Manning dozat_deep_2017 , entails modeling word interactions as a fully connected graph via biaffine attention. Despite its simplicity, a number of models using this biaffine transformation bhatt_end–end_2024 ; dozat_deep_2017 ; dozat_simpler_2018 ; ji_graph-based_2019 ; jiang_entity-relation_2024 require more parameters than necessary, due to what we identify as a lack of normalization of its outputs. We hypothesize that this overparameterization is caused by the high variance of the outputs, which extra parameters help mitigate. After showing that variance can be reduced through normalization, we demonstrate that similar or better performance can be obtained for DP with fewer parameters and training samples.
In this task, we consider a graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$ comprising a set of nodes $\mathcal{V}$, connected by a set of edges $\mathcal{E}$, with each edge $e_{ij}\in\mathcal{E}$ connecting pairs of nodes $(v_{i},v_{j})\in\mathcal{V}$. In latent graph inference (LGI) for dependency graphs, a sentence of $|\mathcal{V}|$ words is modeled as a $|\mathcal{V}|\times|\mathcal{V}|$ graph via biaffine scoring dozat_deep_2017 :
@@
Observe that this is equivalent to the unnormalized scores used in vaswani_attention_2017 ’s self-attention: $QK^{\top}/\sqrt{d_{k}}\\$ where $\text{Attention}(Q,K,V)=\text{Softmax}\left(QK^{\top}/\sqrt{d_{k}}\right)V$ and $d_{k}$ is the output size of the keys.
The main difference resides in the fact that Transformer attention scores are scaled by $a=1/\sqrt{d_{k}}$. As explained in vaswani_attention_2017 , this scaling is done because high variance inputs to the softmax function will result in large values dominating the output ($e^{x\gg 1}$) and small values decaying ($e^{x\ll 1}$). The downstream effect is exploding and vanishing gradients. To see how this works, observe that the score $s$ between any query and key vector is the result of their dot product $q_{i}k_{i}^{\top}$. If we assume that these vectors are zero mean and unit variance, then the variance is $\text{Var}(s)=d_{k}$. Finally, we get our normalization factor, which ensures that each entry in the score matrix has a standard deviation of 1, since $\text{Std}=\sqrt{d_{k}}$. Consequently, lower input variance will result in more stable outputs.
We observe that there is no consistency in the literature on the use of this scaling term for DP. Most works do not use this scaling term  bhatt_end–end_2024 ; dozat_deep_2017 ; dozat_simpler_2018 ; ji_graph-based_2019 ; jiang_entity-relation_2024 , with the exception of tang-etal-2022-unirel who use vaswani_attention_2017 ’s self-attention out of the box. In this paper, we carry out a series of experiments on semantic and syntactic DP tasks, along with non-linguistic LGI tasks, using a wide range of architecture ablations. We extend the work of Bhatt et al. bhatt_end–end_2024 , which represents the state of the art on dependency graph parsing in NLP.
@@
The rest of the paper is structured as follows. Section 2 presents an overview of the literature concerning DP, with specific focus on how to infer the adjacency matrix of a dependency graph. Section 3 explains why layer depth can compensate for normalization. Section 4 provides an overview of the datasets, models, evaluation, and other experimental details. Section 5 illustrates how normalization mitigates overparameterization. Finally, Section 6 draws conclusions and provides avenues for future research.
@@ 2 Background
LGI involves predicting all, or a proper subset of, the edges between the nodes of a graph kazi_differentiable_2023 ; lu2023latent ; velickovic_pointer_2020 . For example, one could carry out LGI over molecule graphs ramakrishnan2014qm9 to predict the bonds between atoms or the proximity of pixel clusters in an image dwivedi2023benchmarking .
In the context of NLP, DP is a task concerned with inferring the dependency graph of a sentence qi2019universal . Depending on the nature of the nodes and relations of the graph, it can be described as semantic (SemDP) bhatt_end–end_2024 ; dozat_simpler_2018 ; jiang_entity-relation_2024 ; tang-etal-2022-unirel or syntactic (SynDP) dozat_deep_2017 ; ji_graph-based_2019 ; yang2018scidtb . In this work, we tackle both tasks, but focus more specifically on SemDP, which can be thought of as comprising two sub-tasks: named entity recognition (NER) nadeau_survey_2009 and relation extraction (RE) bhatt_end–end_2024 ; jiang_entity-relation_2024 . They can be either approached in a pipeline as separate objectives chan2011exploiting ; yan-etal-2021-partition ; zelenko2003kernel ; zhong-chen-2021-frustratingly or by jointly training a single model on both sub-tasks bhatt_end–end_2024 ; bi_codekgc_2024 ; jiang_entity-relation_2024 ; zaratiana_autoregressive_2024 . In the context of deep learning models trained end-to-end, three main paradigms are used: encoder-based models bhatt_end–end_2024 ; dozat_deep_2017 ; dozat_simpler_2018 ; ji_graph-based_2019 ; jiang_entity-relation_2024 ; wadden-etal-2019-entity ; wang-lu-2020-two , decoder-only Transformers, more commonly known as large language models (LLMs) bi_codekgc_2024 ; wei_chatie_2024 ; zhang_extract_2024 , and seq2seq encoder-decoder Transformers liu-etal-2022-autoregressive ; lu-etal-2022-unified ; paolini2021structured ; zaratiana_autoregressive_2024 .
In this work, we focus on encoder-style models, since they currently achieve the best performance on DP tasks bhatt_end–end_2024 ; jiang_entity-relation_2024 ; tang-etal-2022-unirel and are much more parameter efficient than LLM-based solutions. These models approach entity (node) prediction analogously to NER, while edges and relations are handled via MLP projection of node-pair features kiperwasser_simple_2016 ; pei_effective_2015 ; peng-etal-2017-deep ; zhu-etal-2019-graph or via attention-based edge and relation inference bhatt_end–end_2024 ; dozat_simpler_2018 ; ji_graph-based_2019 ; jiang_entity-relation_2024 ; tang-etal-2022-unirel . Current state-of-the-art models, exemplified by bhatt_end–end_2024 , are based on the biaffine dependency parser introduced by dozat_deep_2017 , which uses stacked BiLSTM layers on top of embeddings from a pre-trained language model.
@@ 3 Layer depth can compensate for normalization
We explain this discrepancy using insights from the theory of implicit regularization arora_implicit_2019 , which contains the result that, as layer depth increases, the effective rank of the weight matrices is reduced during gradient descent. Our claim is that a reduction in the rank of the weight matrices causes a corresponding decrease in the variance of the input to the softmax function.
@@
Minimization of a loss function $\mathcal{L}(W)$ with gradient descent using weight matrices $\mathbf{W}$ (assuming a small learning rate $\eta$ and initialization close to the origin). An $N$-layer linear neural network leads the singular values $\sigma_{r}$ of  $\mathbf{W}$ to evolve in the number of iterations $t$ by:
$\sigma_{r}(t+1)\leftarrow\sigma_{r}(t)-\eta\cdot\langle\nabla\mathcal{L}(W(t)),\mathbf{u}_{r}(t)\mathbf{v}_{r}^{\top}(t)\rangle\cdot N\cdot\sigma_{r}(t)^{2-2/N},$
where $\mathbf{u}_{r}\mathbf{v}_{r}$ are the corresponding right and left singular vectors. This implies that, as the number of layers increases, the rank of the weight matrices decreases, since the smallest singular values decay.
Figure 1: Effective rank $\rho(W)$ reduces over training epochs as we increase NN of BiLSTM layers.
While Result 1 is stated for deep linear neural networks for the matrix factorization task, it has been validated empirically as applying to deep non-linear neural networks razin2020implicit ; zhao2022combining . We validate this finding for the biaffine scoring setting in Figure 1, which shows the inverse proportionality between the effective rank $\rho(W)$ roy2007effective of the weights of a BiLSTM and the number $N$ of its layers.
Claim 1 (Monotonic Increase of Output Variance with Rank).
Let $X\in\mathbb{R}^{n}$ be a random vector with covariance matrix $\mathbf{K}_{xx}:=\text{Cov}(X)\in\mathbb{R}^{n\times n}$, and let $\mathbf{A}\in\mathbb{R}^{m\times n}$ be a fixed matrix with singular value decomposition (SVD):
$\mathbf{A}=\sum_{i=1}^{min(m,n)}\sigma_{i}\mathbf{u}_{i}\mathbf{v}_{i}^{\top},$
where $\sigma_{1}\geq\sigma_{2}\geq...\geq 0$. Define the best rank-$r$ approximation of $\mathbf{A}$ by truncated SVD as:
$\mathbf{A}_{r}:=\sum_{i=1}^{r}\sigma_{i}\mathbf{u}_{i}\mathbf{v}_{i}^{\top}.$
Let $Y_{r}=\mathbf{A}_{r}X\in\mathbb{R}^{m}$ denote the image of $X$ under the rank-$r$ linear map. Then, the total variance of $Y_{r}$, as measured by the trace of its covariance matrix, increases monotonically with $r$:
$\text{tr}(\text{Cov}(Y_{r}))\leq\text{tr}(\text{Cov}(Y_{r+1})).$
@@
Let $\mathbf{K}_{xx}:=\text{Cov}(X)\in\mathbb{R}^{n\times n}$ denote the covariance matrix of the random vector $X$. Since covariance matrices are symmetric and positive semi-definite (PSD), we have $\mathbf{K}_{xx}\succeq 0$. Let $\mathbf{A}_{r}:=\sum_{i=1}^{r}\sigma_{i}\mathbf{u}_{i}\mathbf{v}_{i}^{\top}$ be the rank-$r$ truncated SVD of $\mathbf{A}$, the covariance matrix of $Y_{r}:=\mathbf{A}_{r}X$ is:
$\text{Cov}(Y_{r})=\mathbf{A}_{r}\mathbf{K}_{xx}\mathbf{A}_{r}^{\top}.$
To evaluate the total variance we compute the trace:
$\text{tr}\left(\text{Cov}(Y_{r})\right)=\text{tr}\left(\mathbf{A}_{r}\mathbf{K}_{xx}\mathbf{A}_{r}^{\top}\right).$
Using the cyclic property of the trace $\left(\text{tr}(ABC)=\text{tr}(BCA)\right)$ and symmetry of $\mathbf{K}_{xx}$ we get:
$\text{tr}\left(\mathbf{A}_{r}\mathbf{K}_{xx}\mathbf{A}_{r}^{\top}\right)=\text{tr}\left(\mathbf{K}_{xx}\mathbf{A}_{r}^{\top}\mathbf{A}_{r}\right).$
Now define the matrix $\mathbf{M}_{r}:=\mathbf{A}_{r}^{\top}\mathbf{A}_{r}\in\mathbb{R}^{n\times n}$, which is PSD. As $r$ increases, $\mathbf{A}_{r}$ includes more terms in its truncated SVD, so we have:
$\mathbf{M}_{r}=\sum_{i=1}^{r}\sigma^{2}_{i}\mathbf{v}_{i}\mathbf{v}_{i}^{\top},\quad\mathbf{M}_{r+1}=\mathbf{M}_{r}+\sigma^{2}_{r+1}\mathbf{v}_{r+1}\mathbf{v}_{r+1}^{\top}.$
Thus:
$\mathbf{M}_{r+1}\succeq\mathbf{M}_{r}.$
Since $\mathbf{K}_{xx}\succeq 0$, and since the trace of a product of PSD matrices respects Loewner order (i.e., if $A\preceq B$, then $\text{tr}(CA)\leq\text{tr}(CB)$ for all $C\succeq 0$), we conclude:
$\text{tr}(\mathbf{K}_{xx}\mathbf{M}_{r})\leq\text{tr}(\mathbf{K}_{xx}\mathbf{M}_{r+1}),$
which implies:
$\text{tr}\left(\text{Cov}(Y_{r})\right)\leq\text{tr}\left(\text{Cov}(Y_{r+1})\right).$
Therefore, as the rank $r$ increases, the total variance $Y_{r}$ increases. ∎